# Recovering Injected Astrophysics from the LISA Galactic Binaries

### v1

@@ 4.2 Characterizing the likelihood of individual LISA-bright sources
Scott’s rule has been used historically to construct Kernel Density Estimates (KDEs) from a set of samples, where a multivariate normal distribution is constructed to estimate the contribution of each sample to the probability density. For now, we are not truly constructing a KDE, as we evaluate the likelihood on a source-by-source basis, our number of samples is determined by our galactic population, the correlation between parameters is preserved, and because Scott’s rule only overrides the covariance estimates for which a smaller covariance was predicted by the Fisher information. Our likelihood model is a Gaussian Mixture Model (GMM), in the sense that we are sensitive to individual modes of the population (one for each source).
@@ 4.3 A one-dimensional recovery of $\alpha_{\mathrm{CE}}$ from an injected population
For the LISA-bright DWD binaries, we’re not recording a number of instantaneous events in time, but we are still interested in counting the total number of sources identified in an injected population. In this work, we assume that the Inhomogeneous Poisson Point Process likelihood can be used to measure the agreement of an injected population of LISA-bright DWD binaries with test populations. We use the Inhomogeneous Poisson Point Process likelihood (Eq. 12). to evaluate the agreement of our injected sources with test populations.
@@ 5 Constructing a KL divergence statistic for a population
In order to guarantee that we are constructing a minimally biased model, it is useful to use more than one statistic to characterize the agreement of our injected population with test populations. In addition to the Inhomogeneous Poisson Point Process likelihood, we also use the Kullback-Leibler (KL) divergence (Kullback, 1959) between two populations. The KL divergence between two distributions $p,q$ is defined by the asymmetric expression
@@
Critically, the KL divergence is non-negative-definite, and can be zero if and only if $p=q$. In our case, for any two populations, we can make predictions of their observed properties (both in expected number $\mu$ and distribution $p$ in $f_{\rm GW},{\cal M}_{c}$), following the same framework described above, and then compare how similar their predictions should be, either to one another or to a third reference distribution. Following O’Shaughnessy (2013), these differences are intimately related to the (expected) likelihood difference between these predictions and the true model parameters.
In the following discussion, we will fix the injected population, then generate a new bootstrapped population model based on the observed sample. This bootstrapped population will predict on average $\mu_{*}=N$ observations, with a distribution $p_{*}$ of WD binary parameters $f_{\rm GW},{\cal M}_{c}$ estimated by the same likelihood-based method described above, treating treating our set of source likelihoods as a KDE with equally weighted components. In the construction of the KDE, we found that using a further inflated covariance estimate to define the width of each Gaussian component enabled a more informative recovery in $\alpha_{\mathrm{CE}}$. The covariance of each component is expanded to twice Scott’s rule. The maximum contribution of a single sample is also limited by setting $\ln p/q\leq 1$ for individual samples.
@@
Figure 7: The KL divergence in $\alpha_{\mathrm{CE}}$: Taking the same view as Figure 6, we see the components of the KL divergence attributed to Poisson error (top), the source distribution (middle), and both of these together (bottom).
@@
As a concrete example, for our 1-parameter model, Figure 7 shows the KL divergence between our bootstrapped population and models with different $\alpha_{\rm CE}$. In the top panel, we show the KL divergence calculated using only information about the expected number of events. Specifically this panel is obtained by computing the following KL divergence between two Poisson distributions with expected event counts $\mu_{*},\mu$
$D_{\rm KL}=\mu-\mu_{*}+\mu_{*}\ln(\mu_{*}/\mu)\;.$ (18)
In the second panel, we compute the KL divergence derived using only the WD orbital parameter distributions $p(x)$. In the third panel, we show an overall combination:
$D_{\rm KL,joint}=D_{\rm KL}(\mu_{*}|\mu)+\mu_{*}D_{\rm KL}(p_{*}|p).$ (19)
In all three cases, the minimum value of the KL divergence occurs (as expected) at the true parameter value, even allowing for different galaxy realizations.
@@ 5.1 A relationship between likelihood and KL divergence
To understand why the KL divergence and likelihood provide similar information, we review their expected relationship. Any proposed realization of our bootstrapped population produces some number $n$ of events, with parameters $x_{k}$ each independently drawn from the bootstrapped distribution $p(x)$. Here, $x$ are phenomenological parameters of the population (here $f_{\rm GW}$ and ${\cal M}_{c}$). We can compare any realization to a proposed population model, with expected number of events $\mu$ and phenomenological population $p$, with the standard inhomogeneous Poisson likelihood $P(n|\mu)\prod_{k}p(x_{k})$. Averaging over different realizations of the (bootstrapped) source population model, we find
@@
where we note ${\left\langle n\right\rangle}=\mu_{*}$. Following O’Shaughnessy (2013) and reorganizing, we can express the difference in (expected) log likelihood between any two candidate populations $(\mu_{A},p_{A})$ and $(\mu_{B},p_{B})$ characterized by $\Lambda_{\mathrm{A}}$ and $\Lambda_{\mathrm{B}}$ in terms of differences in KL divergences:
@@
$\mu_{*}[\mathrm{D}_{\mathrm{KL}}(p|p_{A})-\mathrm{D}_{\mathrm{KL}}(p|p_{B})]$
In the special case that $B$ is the (bootstrapped) reference population, then the likelihood difference will be zero if and only if $A$ is also the bootstrapped reference population, and both $D_{KL}$ terms contributing to the likelihood are nonnegative-definite. Figure 8 validates this relationship, showing the close agreement between the rate likelihood and the number-only KL divergence $D_{KL}(\mu_{*}|\mu)$; the shape likelihood and the shape-only KL divergence $\mu_{*}D_{\rm KL}(p_{*}|p)$ ; and the overall likelihood and the “joint” KL divergence introduced above.
Figure 8: Agreement of the KL divergence and log likelihood: Taking the same view again as Figure 6, we measure the agreement of two methods of constraint: the log likelihood and KL divergence. These values should all be close to 1.0 on the y axis.
@@ 7 Discussion
In this work we have assessed the impact of common envelope evolution uncertainties on the population of DWDs that LISA may discover. We simulated 480 populations of DWDs across 15 metallicities and 32 common envelope ejection efficiencies using COSMIC. We then applied the metallicity-specific star formation history of the Milky-Way like galaxy m12i from the FIRE-2 Latte Simulation Suite to create a synthetic Milky-Way-like population of DWDs. With this synthetic population, we investigated LISA’s ability to distinguish between models for the common envelope ejection efficiency using the measured strain and gravitational wave frequency evolution. Finally, we explored the potential for an inhomogeneous Poisson point process likelihood to make those distinctions and compared with predictions made using a KL divergence approach, through both the number and properties of predicted detections.
@@
  • The common envelope ejection efficiency, $\alpha_{\mathrm{CE}}{}$, directly correlates with the number of DWDs that form per unit solar mass. This is due to fewer pre-DWD formation stellar mergers in failed common envelopes with increasing ejection efficiency.
@@
  • These trends in DWD binary formation impact the size and properties of the galactic population of DWDs, but are reduced in the population of resolved LISA DWDs. This is due to the interplay of the DWD confusion foreground, which may reduce the number of resolved sources for larger galactic populations.
  • Both the inhomogeneous Poisson point process likelihood and KL divergence statistics can be used to quantitatively understand the distinction between models of binary formation.
  • The strongest distinctions are found when constraining formation models with both the number and properties of LISA observable DWD systems. The predicted number of detections alone may not be sufficient to distinguish between models.
  • Our study predicts that LISA may be sensitive to changes in common envelope ejection efficiency at about the $10\%$ level. With the caveat that we did not account for the uncertainty associated with the stability of mass transfer which determines the fraction of initial binaries that undergo common envelope evolution, our findings suggest that LISA observations may play a key role in understanding binary evolution. Specifically, phases of mass transfer initiated by Roche-lobe overflow may have an observable impact on the population of LISA DWD binaries if the dominant channel that produces them is common envelope evolution.
@@
The authors thank Ann E. Hornschemeier for useful feedback. VD is supported by an appointment to the NASA Postdoctoral Program at the NASA Goddard Space Flight Center administered by Oak Ridge Associated Universities under contract NPP-GSFC-NOV21-0031. ROS gratefully acknowledges support from NSF awards NSF PHY-1912632, PHY-2012057, PHY-2309172, AST-2206321, and the Simons Foundation. JB is supported by the NASA LISA Project Office. We acknowledge software packages used in this publication, including NUMPY (Harris et al., 2020), SCIPY (Virtanen et al., 2020), MATPLOTLIB (Hunter, 2007), ASTROPY (Astropy Collaboration et al., 2013, 2018), H5PY (Collette, 2013), LEGWORK (Wagg et al., 2022), and ldasoft Littenberg et al. (2020). This research was done using resources provided by the Open Science Grid (Pordes et al., 2007; Sfiligoi et al., 2009), which is supported by the National Science Foundation awards #2030508 and #1836650, and the U.S. Department of Energy’s Office of Science.

### v2

@@ 4.2 Characterizing the likelihood of individual LISA-bright sources
Scott’s rule has been used historically to construct Kernel Density Estimates (KDEs) from a set of samples, where a multivariate normal distribution is constructed to estimate the contribution of each sample to the probability density. For now, we are not truly constructing a KDE, as we evaluate the likelihood on a source-by-source basis, our number of samples is determined by our Galactic population, the correlation between parameters is preserved, and because Scott’s rule only overrides the covariance estimates for which a smaller covariance was predicted by the Fisher information. Our likelihood model is a Gaussian Mixture Model (GMM), in the sense that it is sensitive to individual modes of the population (one for each source).
@@ 4.3 A one-dimensional recovery of $\alpha_{\mathrm{CE}}$ from an injected population
For the LISA-bright DWD binaries, we’re not recording a number of instantaneous events in time, but we are still interested in counting the total number of sources identified in an injected population. In this work, we assume that the Inhomogeneous Poisson Point Process likelihood can be used to measure the agreement of an injected population of LISA-bright DWD binaries with test populations. We use the Inhomogeneous Poisson Point Process likelihood (Eq. 13) to evaluate the agreement of our injected sources with test populations.
@@ 5 Constructing a KL divergence statistic for a population
In order to guarantee that we are constructing a minimally biased model, it is useful to use more than one statistic to characterize the agreement of our injected population with test populations. In addition to the Inhomogeneous Poisson Point Process likelihood, we also use the Kullback-Leibler (KL) divergence (Kullback, 1959) between two populations. We use the KL divergence to measure the information lost by assuming that a set of observations ($p_{*}$) can describe a modeled population ($p$). Similar to Section 4, we describe a particular population by a number of observations ($\mu$ or $\mu_{*}$) and a discrete set of samples which represent a density $p(\vec{\lambda})$ or $p_{*}(\vec{\lambda})$. For the KL divergence approach, we characterize the density function with the $\log_{10}(f_{\mathrm{GW}})$ and $\mathcal{M}_{c}$ binary parameters ($\vec{\lambda}$).
5.1 The Kullback-Leibler divergence
Similar to the Inhomogeneous Poisson Point Process likelihood, the KL divergence between a tested population and reference population can be estimated with information about the rate of each population ($D_{\mathrm{KL}}(\mu_{*}|\mu)$, for a number of observations $\mu$), shape of each population ($D_{\mathrm{KL}}(p_{*}|p)$), or both ($D_{\mathrm{KL}}(p_{*},\mu_{*}|p,\mu)$). We test the compatibility of a set of injected populations ($p_{*}$) with several modeled populations ($p$).
The KL divergence between our injected observations ($p_{*}$) and a particular test model ($p$) is defined by the asymmetric expression:
@@ 5.1 The Kullback-Leibler divergence
The KL divergence between a number of injected or observed events ($\mu_{*}$) and an expected number of events ($\mu$) can be derived from the Poisson distribution, and is simple to evaluate numerically:
$D_{\mathrm{KL,rate}}=D_{\mathrm{KL}}(\mu_{*}|\mu)\equiv\mu_{*}-\mu+\mu\ln(\mu/\mu_{*})\;.$ (19)
A measure of KL divergence informed by both rate and shape information must scale accordingly with $\mu_{*}$:
$D_{\mathrm{KL,joint}}=D_{\mathrm{KL}}(p,\mu|p_{*},\mu_{*})\equiv D_{\rm KL}(\mu_{*}|\mu)+\mu_{*}D_{\rm KL}(p_{*}|p).$ (20)
The KL divergence is positive semi-definite (it will be zero if and only if $p=p_{*}$).
By constructing a functional form to represent the density of each population model, and integrating over the discrete set of samples which define observed population ($p_{*}$), we can define a discretized KL divergence:
$D_{\mathrm{KL,shape}}=\frac{1}{\mu_{*}}\sum\limits_{k}\ln(p_{*}(\vec{\lambda}_{k}))-\ln(p(\vec{\lambda}_{k}))$ (21)
where $p(\vec{\lambda}_{k})$ and $p_{*}(\vec{\lambda}_{k})$ are approximations to the density of sample binaries which make up $p$ and $p_{*}$ respectively (The construction of which is deferred to Section 5.2). This is a Monte Carlo integration over samples from $p_{*}$ (indexed by $k$), and this sampling accounts for the factor of $p_{*}$ outside the logarithm in Eq. 18.
5.2 Characterizing the Density of LISA-bright sources
When carrying out the Monte Carlo integration in Eq. 21, we must construct an approximation to the sample densities that make up $p$ and $p_{*}$. These approximations take the form of a Gaussian mixture model, where $p(\vec{\lambda})=\sum_{j}\mathcal{L}_{j}(\vec{\lambda})/\mu$ and $p_{*}(\vec{\lambda})=\sum_{j*}\mathcal{L}_{j*}(\vec{\lambda})/\mu_{*}$. Here, $\mathcal{L}_{j}$ and $\mathcal{L}_{j*}$ are individual Gaussian components of the mixture model, and we construct them from the PE likelihood of each LISA-bright source (i.e. 12). We make the same SNR and $f_{\mathrm{GW}}$ cuts as in Section 4, and construct each population with the LISA-bright sample binaries. However, we make a modification by changing coordinates from $f_{\mathrm{GW}}$ to $\log(f_{\mathrm{GW}})$ in the construction of our density estimate. We also further inflate covariance estimates to define the width of each Gaussian component by a minimum of twice Scott’s rule Scott (2015).
When sampling many galaxy realizations for each $\alpha_{\mathrm{CE}}$ value, we must choose a width for our Gaussian components which is wide enough so that the sampling error between realizations (sampled with different random number generator seeds) are not greater than the difference between our assumptions for $\alpha_{\mathrm{CE}}$. While it may be possible to construct a mathematical method for choosing the width of Gaussian components to overcome this sampling error, we chose this value by hand in this work. The impact of individual samples from the test population on the KL divergence is also limited by setting a maximum contribution ($\ln p/p_{*}\leq 1$) for individual samples to the integration in Eq. 21. As both the Inhomogeneous Poisson Point Process likelihood and KL divergence models require inflated covariances to at least Scott’s rule, a detailed parameter estimation likelihood model may be unnecessary. As the KL divergence is more susceptible to sampling error, a likelihood-based approach may be preferable.
@@ 5.2 Characterizing the Density of LISA-bright sources
Figure 7: The KL divergence in $\alpha_{\mathrm{CE}}$: Taking the same view as Figure 6, we see the components of the KL divergence attributed to Poisson error (top), the shape of the source distribution (middle), and both of these together (bottom). We note the presence of a few realizations for which the estimated KL divergence appears to be numerically unstable, despite our choices of coordinates and inflated Gaussian components.
@@
As a concrete example of this KL divergence approach Figure 7 shows the KL divergence between our Gaussian mixture models representing the injected CEb60 population and test populations with different $\alpha_{\mathrm{CE}}$. In the top panel, we show the KL divergence calculated using only information about the expected number of events. In the second panel, we compute the KL divergence derived using only the DWD orbital parameter distributions $p(\vec{\lambda})$. In the third panel, we show an overall combination: In all three cases, the minimum value of the KL divergence occurs (as expected) near the true parameter value, even allowing for different galaxy realizations.
@@ 5.3 A relationship between likelihood and KL divergence
To understand why the KL divergence and likelihood provide similar information, we review their expected relationship. This section is heavily inspired by O’Shaughnessy (2013).
Just as we are interested in finding the more likely model when using a maximum likelihood approach (rather than fixating on the value of a particular likelihood estimate), we are more interested here in the difference between KL divergence values for particular models (rather than the value of a particular KL divergence estimate). Any proposed realization of a “bootstrapped” reference population that approximates the samples belonging to $p_{*}$ will predict on average $\mu_{*}$ observations, with a distribution $p_{*}$, consistent with the injected population model. Following O’Shaughnessy (2013), the difference in KL divergence between the bootstrapped population and a given test population (characterized by $\mu$ observations and a distribution of $p$) are intimately related to the (expected) likelihood difference between the tested population model and injected population model. Averaging over possible realizations of the bootstrapped source population model, we find
@@
where we note ${\left\langle n\right\rangle}=\mu_{*}$.
Reorganizing, we can express the difference in (expected) log likelihood between any two candidate populations $(\mu_{A},p_{A})$ and $(\mu_{B},p_{B})$ characterized by $\Lambda_{\mathrm{A}}$ and $\Lambda_{\mathrm{B}}$ in terms of differences in KL divergences:
@@
$\mu_{*}[\mathrm{D}_{\mathrm{KL}}(p_{*}|p_{A})-\mathrm{D}_{\mathrm{KL}}(p_{*}|p_{B})]$
In the special case that $B$ is our mixture model approximating the (injected) reference population, then the likelihood difference will be zero if and only if $A$ is also the bootstrapped reference population. Critically, we can therefore check if an estimated KL divergence and likelihood estimate agree with our understanding of their construction by dividing the left half of this equation by the right. If our model is well-constructed, the resulting expression should be consistent with unity.
Figure 8 validates this relationship, showing the close agreement between the rate likelihood and the Poisson KL divergence $D_{\mathrm{KL,rate}}(\mu_{*}|\mu)$; the shape likelihood and the shape-only KL divergence $\mu_{*}D_{\mathrm{KL,shape}}(p_{*}|p)$ ; and the overall likelihood and the “joint” KL divergence introduced above.
The key differences between this approach and the Inhomogeneous Poisson Point Process likelihood are:
  • The KL divergence is sensitive to a difference between two models, rather than the agreement of a single model to a fixed set of observations (compare Eq. 18 and Eq. 13).
  • In the likelihood approach, we only needed to construct a Gaussian mixture model for the injected population (not each test population). In the KL divergence approach, $p$ and $p_{*}$ are both evaluated using the functional form approximation (i.e. the bootstrapped model) from their own individual population samples.
Figure 8: Agreement of the KL divergence and log likelihood: Taking the same view again as Figure 6, we measure the agreement of two methods of constraint: the log likelihood and KL divergence. As per Eq. 23, these values should be close to 1.0 if our model is well-constructed. We note that this is essentially true in the vicinity of the injected model ($\alpha_{\mathrm{CE}}=1.0$).
@@ 7 Discussion
In this work we have assessed the impact of common envelope evolution uncertainties on the population of DWDs that LISA may discover. We simulated 480 populations of DWDs across 15 metallicities and 32 common envelope ejection efficiencies using COSMIC. We then applied the metallicity-specific star formation history of The Milky-Way-like galaxy m12i from the FIRE-2 Latte Simulation Suite to create a synthetic Milky-Way-like population of DWDs. With this synthetic population, we investigated LISA’s ability to distinguish between models for the common envelope ejection efficiency using the measured strain and gravitational wave frequency evolution. Finally, we explored the potential for an Inhomogeneous Poisson point process likelihood to make those distinctions and compared with predictions made using a KL divergence approach, through both the number and properties of predicted detections.
@@
  • The common envelope ejection efficiency, $\alpha_{\mathrm{CE}}{}$, directly correlates with the number of DWDs that form per unit solar mass. As $\alpha_{\mathrm{CE}}$ increases, the number of stellar mergers from failed common envelope ejections among the progenitors of DWD binaries decreases.
@@
  • The effects of changing $\alpha_{\mathrm{CE}}$ mentioned above are also seen in the LISA Galactic DWD and LISA bright populations; however, the LISA bright population is less effected due to competition with the DWD confusion foreground.
  • Both the Inhomogeneous Poisson point process likelihood and KL divergence statistics can be used to quantitatively distinguish between different models for DWD formation. We find that because the KL divergence method is more susceptible to sampling error, we reccommend the Inhomogenous Poisson Point Process likelihood for future population studies.
  • Because inflated Gaussian covariances which are larger than PE uncertainties are required to overcome sampling error, we find that applying Scott’s rule makes the Fisher estimates unnecessary.
  • The strongest Bayesian inference constraints are found when constraining formation models with both the number and properties of LISA observable DWD systems. The predicted number of detections alone may not be sufficient to distinguish between models.
  • When considering only changes to common envelope ejection efficiency, we find that LISA is sensitive at about the 10 % level. Future studies should consider how constraints change when incorporating other uncertain models (E.g. different common envelope models, mass transfer stability).
@@
Data/Code availability
A limited set of data supporting the findings of this study is openly available at https://zenodo.org/records/13872540(Delfavero et al., 2024a). The code supporting the findings of this study is available at https://zenodo.org/records/14421368(Delfavero et al., 2024b).
The authors thank Ann E. Hornschemeier and the anonymous referee for useful proofreading and feedback. VD is supported by an appointment to the NASA Postdoctoral Program at the NASA Goddard Space Flight Center administered by Oak Ridge Associated Universities under contract NPP-GSFC-NOV21-0031. ROS gratefully acknowledges support from NSF awards NSF PHY-1912632, PHY-2012057, PHY-2309172, AST-2206321, and the Simons Foundation. JB is supported by the NASA LISA Project Office. We acknowledge software packages used in this publication, including NUMPY (Harris et al., 2020), SCIPY (Virtanen et al., 2020), MATPLOTLIB (Hunter, 2007), ASTROPY (Astropy Collaboration et al., 2013, 2018; ast, 2022), H5PY (Collette, 2013), LEGWORK (Wagg et al., 2022), and ldasoft Littenberg et al. (2020). This research was done using resources provided by the Open Science Grid (Pordes et al., 2007; Sfiligoi et al., 2009), which is supported by the National Science Foundation awards #2030508 and #1836650, and the U.S. Department of Energy’s Office of Science.